hamnakahlid:

mixture placed in an Erlenmeyer flask comprises 6mL of silica gel and 40mL of a solvent containing, in solution, 100 mg of a non-volatile compound. After stirring, the mixture was left to stand before a 10mL aliquot of the solution was extracted and evaporated to dryness. The residue weighed 12 mg. Calculate the adsorption coefficient, K = CS/CM, of the compound in this experiment

adrianluvvsyouu2:

Start by calculating Cm, which is Mass in aliquot divided by its volume. So you said that the aliquot was 10mL (volume) and its residue weighed 12mg... 12/10 = 1.2 mg/mL which is your Cm To find the total mass of our solution that remains, we can do Cm * Total volume (since the units cancel out) 1.2 mg/mL * 40mL = 48mg of solution remaining Since we started with 100mg and are left with 48mg, we know something must have happened to the 52mg, it was absorbed. We can use this to find the value of Cs, the concentration of the compound in the silica gel, which has a volume of 6 mL 52mg absorbed/6 mL volume of silica gel = 8.67 mg/mL = Cs Now that we have Cm and Cs, we can do Cs/Cm which is our K absorption coefficient... 8.67 mg/mL divided by 1.2 mg/mL, the units cancel and we get a coefficient of 7.23.

ihy:

@adrianluvvsyouu2 wrote:
Start by calculating Cm, which is Mass in aliquot divided by its volume. So you said that the aliquot was 10mL (volume) and its residue weighed 12mg... 12/10 = 1.2 mg/mL which is your Cm To find the total mass of our solution that remains, we can do Cm * Total volume (since the units cancel out) 1.2 mg/mL * 40mL = 48mg of solution remaining Since we started with 100mg and are left with 48mg, we know something must have happened to the 52mg, it was absorbed. We can use this to find the value of Cs, the concentration of the compound in the silica gel, which has a volume of 6 mL 52mg absorbed/6 mL volume of silica gel = 8.67 mg/mL = Cs Now that we have Cm and Cs, we can do Cs/Cm which is our K absorption coefficient... 8.67 mg/mL divided by 1.2 mg/mL, the units cancel and we get a coefficient of 7.23.
Ai?

adrianluvvsyouu2:

@ihy wrote:
@adrianluvvsyouu2 wrote:
Start by calculating Cm, which is Mass in aliquot divided by its volume. So you said that the aliquot was 10mL (volume) and its residue weighed 12mg... 12/10 = 1.2 mg/mL which is your Cm To find the total mass of our solution that remains, we can do Cm * Total volume (since the units cancel out) 1.2 mg/mL * 40mL = 48mg of solution remaining Since we started with 100mg and are left with 48mg, we know something must have happened to the 52mg, it was absorbed. We can use this to find the value of Cs, the concentration of the compound in the silica gel, which has a volume of 6 mL 52mg absorbed/6 mL volume of silica gel = 8.67 mg/mL = Cs Now that we have Cm and Cs, we can do Cs/Cm which is our K absorption coefficient... 8.67 mg/mL divided by 1.2 mg/mL, the units cancel and we get a coefficient of 7.23.
Ai?
No I took chemistry?

ihy:

@adrianluvvsyouu2 wrote:
@ihy wrote:
Oh, my fault.
@adrianluvvsyouu2 wrote:
Start by calculating Cm, which is Mass in aliquot divided by its volume. So you said that the aliquot was 10mL (volume) and its residue weighed 12mg... 12/10 = 1.2 mg/mL which is your Cm To find the total mass of our solution that remains, we can do Cm * Total volume (since the units cancel out) 1.2 mg/mL * 40mL = 48mg of solution remaining Since we started with 100mg and are left with 48mg, we know something must have happened to the 52mg, it was absorbed. We can use this to find the value of Cs, the concentration of the compound in the silica gel, which has a volume of 6 mL 52mg absorbed/6 mL volume of silica gel = 8.67 mg/mL = Cs Now that we have Cm and Cs, we can do Cs/Cm which is our K absorption coefficient... 8.67 mg/mL divided by 1.2 mg/mL, the units cancel and we get a coefficient of 7.23.
Ai?
No I took chemistry?

V4MPZSKII:

w